### v2

@@ 1.2 Related work
A variant of the Gale substitutes property also appeared in Garg et al. (2022a) [^1] The technical report version of Garg et al. (2023a).. This was used to obtain an auction algorithm for the maximum Nash welfare problem with capped SPLC utilities, and the proof uses arguments similar to Section 7. Similar auction algorithms could be developed more generally for Gale substitutes functions.
Fisher markets are a special case of more general Arrow-Debreu (exchange) markets. Computing equilibria in exchange markets is polynomial time for WGS utility functions (see, e.g., Bei et al. (2019a), Codenotti et al. (2005c), Garg et al. (2023a)) and for CES utilities with parameter $-1\leq\rho<0$ Codenotti et al. (2005a), and beyond that the problem is essentially PPAD-hard Chen et al. (2009), even for homogeneous functions, such as Leontief and CES with parameter $\rho<-1$ Chen et al. (2013), Codenotti et al. (2006).
The maximum Nash welfare is a classic objective for allocating goods to agents. It has received significant attention in the literature on social choice and fair division; see, e.g., Moulin (2003). It provides remarkable fairness guarantees also for indivisible goods: under additive valuations, maximum Nash welfare allocations attain the “unreasonable” fairness guarantee of envy-freeness up to one good (EF1) Caragiannis et al. (2019). Furthermore, for more general subadditive valuations, the fairness guarantee of $\tfrac{1}{2}$-approximate envy-freeness up to any good (EFX) can be achieved at a $\tfrac{2}{3}$-approximate maximum Nash welfare allocation Feldman et al. (2024).
@@
The remainder of the paper is organized as follows. Section 2 defines all models and definitions. Section 3 provides a brief overview of the techniques used. Section 4 proves Theorem 1.1, showing the tight bound of $(1/\mathrm{e})^{1/\mathrm{e}}$ on the efficiency of competitive equilibria. Section 5 presents several properties of the Gale demand correspondence and proves Theorem 1.2, showing that Nash welfare maximizing allocations yield $2$-demand approximate-CE. Section 6 proves Theorem 1.3, showing the stronger guarantees of the maximum Nash welfare allocations for Gale-substitutes utility functions. Section 7 proves Theorem 1.4, showing that separable utilities are $\Sigma$-Gale substitutes. Section 8 considers generalized network utility functions and proves Theorems 1.5 and 1.6. Several examples are deferred to the Appendix, and Appendix B presents an omitted proof from Section 8.
2 Preliminaries
Notation
For a function $f\,:\,X\to Y$, let $\operatorname{dom}(f)\subseteq X$ denote its domain. For $x\in\mathbb{R}^{n}$, let $\operatorname{supp}(x)=\{i\,:\,x_{i}\neq 0\}$ denote its support. For $x,y\in\mathbb{R}^{n}$, we use $\left\langle x,y\right\rangle=x^{\top}y$ for the standard scalar product.
@@ 2 Preliminaries
Thus, $\mathbf{g}$ if a supergradient of $f$ if and only if $-\mathbf{g}$ is a subgradient of $-f$. We recall that a function $f$ is convex if and only if $\partial^{\vee}{f}(\mathbf{x})\neq\emptyset$ for every $\mathbf{x}\in\operatorname{dom}(f)$ and concave if and only if $\partial^{\wedge}{f}(\mathbf{x})\neq\emptyset$ for every $\mathbf{x}\in\operatorname{dom}(f)$. For a differentiable convex (concave) function, the unique subgradient (supergradient) at every point is the gradient $\nabla f(\mathbf{x})$.
@@ 2.1 Utility functions
A function $u(\cdot)\,,\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}$ is a utility function if $u(\mathbf{0})=0$, $u(\cdot)$ is concave, monotone non-decreasing, and there exists an $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $u(x)>0$.
We say that a utility function is separable if $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$. We will also distinguish between satiable and non-satiatiable utilities.
@@
A utility function $u(\cdot)$ is non-satiable if and only if, for any $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, one can find $x^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, such that $u(x^{\prime})>u(x)$. Otherwise, we say the utility function is satiable.
@@
${\mathcal{D}}^{u}(\mathbf{p},b)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\arg\max_{x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\left\{u(x)\,:\,\left\langle\mathbf{p},x\right\rangle\leq b\right\}\,.$ (1)
Note that if $u$ is a non-satiable utility function, then $\left\langle\mathbf{p},x\right\rangle=b$ must hold for every $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$.
${\mathcal{D}}^{u}(\mathbf{p},b)$ is always a non-empty compact set when $\mathbf{p}>0$. However, in cases when some prices may be 00, the supremum of utility over the budget constraint might be $\infty$, resulting in unbounded demand. This may happen even when the supremum value is finite, and causes technical difficulties. To circumvent these issues, we adopt the following assumption regarding utility functions $u(\cdot)$:
@@
For any $\mathbf{p}\geq\mathbf{0}$ and $b>0$, if there is a sequence of allocations $\{x^{(k)}\}_{k=1}^{\infty}$ such that $\left\langle\mathbf{p},x^{(k)}\right\rangle$ converges to $b$ and $u(x^{(k)})$ converges to some finite value $M<\infty$, then there exists a finite allocation $x^{*}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $u(x^{*})=M$ and $\left\langle\mathbf{p},x\right\rangle\leq b$.
This assumption always holds if the set of possible allocations is restricted to a compact domain $X\subseteq\mathbb{R}_{\geq 0}^{\mathcal{M}}$; note Nesterov and Shikhman (2018) makes such a compactness assumption. For the unrestricted domain $\mathbb{R}_{\geq 0}^{\mathcal{M}}$, it can also be shown for arbitrary piecewise linear utility functions, in particular, for generalized network utilities defined below.
@@
$\mathcal{GD}^{u}(\mathbf{p},b)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\arg\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle\,.$ (2)
We define the domain of $\mathcal{GD}^{u}$ with respect to prices as
$\operatorname{pdom}(u)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\left\{\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}\leavevmode\nobreak\ :\leavevmode\nobreak\ \exists b>0,\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle<+\infty\right\}\,.$ (3)
Under Assumption 1, $\mathbf{0}\in\operatorname{pdom}(u)$ if and only if $u(.)$ is satiable. Every strictly positive $\mathbf{p}>\mathbf{0}$ is in $\operatorname{pdom}(u)$, and if $u$ is satiable, then $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$ (see Lemma 5.2). In Section 5, we show the following.
@@
Under Assumption 1, for any $\mathbf{p}\in\operatorname{pdom}(u)$ and any $b>0$, $\mathcal{GD}^{u}(\mathbf{p},b)\neq\emptyset$.
Gale demands may behave very differently from the standard demand correspondence. Example A.3 shows the counter-intuitive property increasing the prices $\mathbf{p}$ may actually increase the Gale utility value $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle$.
Recall the notions of bounded linear, separable piecewise linear (SPLC) and Leontief-free (LF) utilities from the Introduction. We now define generalized network utility functions.
@@
First, we introduce the generalized flow model. A generalized flow instance is defined as follows. Let $G=(V,E)$ be a directed graph with a sink node $t\in V$, a supply set $S\subseteq V\setminus\{t\}$, a capacity vector $h\in\mathbb{R}_{\geq 0}^{E}$, and positive gain factors $\gamma\in\mathbb{R}_{>0}^{E}$. We let $\partial^{\mathrm{in}}(S)$ and $\partial^{\mathrm{out}}(S)$ denote the set of arcs entering and leaving a subset $S\subseteq V$, respectively. For a node $v\in V$, we use $\partial^{\mathrm{in}}(v)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\partial^{\mathrm{in}}(\{v\})$ and $\partial^{\mathrm{out}}(v)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\partial^{\mathrm{out}}(\{v\})$. For a vector $f\in\mathbb{R}_{\geq 0}^{E}$, for every arc $e=(v,w)$, we interpret $f_{e}$ as the flow entering $e$ at node $v$. The flow gets multiplied by the gain factor $\gamma_{e}$ while traversing the arc, with $\gamma_{e}f_{e}$ amount reaching node $w$. Thus, the net flow at node $v\in V$ is defined as
$\operatorname{net}_{f}(v)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{e\in\partial^{\mathrm{in}}(v)}\gamma_{e}f_{e}-\sum_{e\in\partial^{\mathrm{out}}(v)}f_{e}\,.$
Given a supply vector $x\in\mathbb{R}_{\geq 0}^{S}$, we say that $f\in\mathbb{R}_{\geq 0}^{E}$ is a feasible generalized flow with supply $x$ and capacity $h$ if $0\leq f\leq h$, and $\operatorname{net}_{f}(v)\geq-x_{v}$ for all $v\in S$, and $\operatorname{net}_{f}(v)\geq 0$ for all $v\in V\setminus(S\cup\{t\})$. The maximum generalized flow value from supply $x$ is the maximum amount of $\operatorname{net}_{f}(t)$ among all feasible flows with supply $x$. It corresponds to the optimum value of the following linear program:
$\max\leavevmode\nobreak$ $\operatorname{net}_{f}(t)$ ($\operatorname{GenflowLP}(x)$)
$\operatorname{net}_{f}(v)$ $\geq-x_{v}\quad\forall v\in S$
$\operatorname{net}_{f}(v)$ $\geq 0\quad\forall v\in V\setminus(S\cup\{t\})$
00 0 $\leq f\leq h\,.$
@@
The function $u:\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ is a generalized network utility function if $u(x)$ is the maximum generalized flow value from supply $x$ in some generalized flow instance where the supply set is $S={\mathcal{M}}$, the maximum flow amount is finite for any supply vector, $u(\mathbf{0})=0$, and there exists $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ with $u(x)>0$.
The following properties are immediate from the definition, and show that these are indeed utility functions in accordance with Definition 2.2.
@@
If $u:\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ is a generalized network utility function, then $u(x)$ is monotone non-decreasing and concave, and satisfies Assumption 1.
@@
An SPLC function can be represented as a generalized network utility function by setting $V={\mathcal{M}}\cup\{t\}$, and for each $j\in{\mathcal{M}}$, adding a capacitated arcs $e=(j,t)$ for each linear segment of $v_{j}$, where $\gamma_{e}$ is the marginal utility on this segment and $h_{e}$ is the length of the segment. Similary, we can represent a Leontief-free utility as a generalized network utility function by setting $V={\mathcal{M}}\cup K\cup\{t\}$; recall $K$ is the set of segments. For each $j\in{\mathcal{M}}$ and $k\in K$, we add an arc $e=(j,k)$ with gain factor $\gamma_{e}=a_{jk}$ and capacity $h_{e}=\infty$. For each segment $k\in K$, we add an arc $e=(k,t)$ with $\gamma_{e}=1$ and $h_{e}=\ell_{k}$.
@@
Assume that agents are competing for access to $m$ machines ${\mathcal{M}}$ that can produce $\ell$ different products. Machine $j$ can produce $a_{jk}$ units of product $k$ per time unit. The revenue of product $k$ is $r_{k}$. There are additionally limits $t_{jk}$ on how much of each product can be produced on each machine by the agent, as well a limit $d_{k}$ on the maximum possible amount producible from $k$. Given $x_{j}$ time on machine $j\in{\mathcal{M}}$, the utility of the agent is the maximum possible revenue subject to these constraints. This can be formulated using a two layer generalized flow network similarly to Leontief-free as above (albeit with additional capacity bounds $t_{jk}$.) The model allows for more complex production networks with possibly multiple layers corresponding to raw materials, intermediary and final products.

### v3

@@ 1.2 Related work
A variant of the Gale substitutes property also appeared in Garg et al. (2022a) [^4] The technical report version of Garg et al. (2023a).. This was used to obtain an auction algorithm for the maximum Nash welfare problem with capped SPLC utilities, and the proof uses arguments similar to those in Appendix C. Similar auction algorithms could be developed more generally for Gale substitutes functions.
Fisher markets are a special case of more general Arrow-Debreu (exchange) markets. Computing equilibria in exchange markets is polynomial time for WGS utility functions (see, e.g., Bei et al. (2019a), Codenotti et al. (2005c), Garg et al. (2023a)) and for CES utilities with parameter $1\leq\rho<0$ Codenotti et al. (2005a), and beyond that the problem is essentially PPAD-hard Chen et al. (2009), even for homogeneous functions, such as Leontief and CES with parameter $\rho<-1$ Chen et al. (2013), Codenotti et al. (2006).
The maximum Nash welfare is a classical objective for allocating goods to agents. It has received significant attention in the literature on social choice and fair division; see, e.g., Moulin (2003). It provides remarkable fairness guarantees also for indivisible goods: under additive valuations, maximum Nash welfare allocations attain the “unreasonable” fairness guarantee of envy-freeness up to one good (EF1) Caragiannis et al. (2019). Furthermore, for more general subadditive valuations, the fairness guarantee of $\tfrac{1}{2}$-approximate envy-freeness up to any good (EFX) can be achieved at a $\tfrac{2}{3}$-approximate maximum Nash welfare allocation Feldman et al. (2024).
@@
The remainder of the paper is organized as follows. Section 2 defines all models and definitions. Section 3 proves Theorem 1.1, establishing the tight bound of $(1/\mathrm{e})^{1/\mathrm{e}}$ on the efficiency of competitive equilibria. In Section 4, we show that Nash welfare maximizing allocations yield $2$-demand approximate-CE. Section 5 proves Theorem 1.3, showing the stronger guarantees of the maximum Nash welfare allocations for Gale-substitutes utility functions. Section 6 proves Theorem 1.4, showing that separable utilities are $\Sigma$-Gale substitutes, and considers generalized network utility functions.
Appendix A presents several properties of the Gale demand correspondence. Appendix B provides the complete proof of Theorem 1.3. Appendix C proves Theorem 1.4 for the general case, and Appendix D contains the proofs of Theorems 1.5 and 1.6. Additional examples are presented in Appendix E.
2 Models and definitions
For a function $f\,:\,X\to Y$, let $\operatorname{dom}(f)\subseteq X$ denote its domain. For $\mathbf{x}\in\mathbb{R}^{n}$, let $\operatorname{supp}(\mathbf{x})=\{i\,:\,x_{i}\neq 0\}$ denote its support. For $\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}$, we use $\left\langle\mathbf{x},\mathbf{y}\right\rangle=\mathbf{x}^{\top}\mathbf{y}$ for the standard scalar product.
@@ 2 Models and definitions
Thus, $\mathbf{g}$ if a supergradient of $f$ if and only if $\mathbf{g}$ is a subgradient of f-f. We recall that a function $f$ is convex if and only if $\partial^{\vee}{f}(\mathbf{x})\neq\emptyset$ for every $\mathbf{x}\in\operatorname{dom}(f)$ and concave if and only if $\partial^{\wedge}{f}(\mathbf{x})\neq\emptyset$ for every $\mathbf{x}\in\operatorname{dom}(f)$. For a differentiable convex (concave) function $f$, the unique subgradient (supergradient) at every point is the gradient $\nabla f(\mathbf{x})$.
@@ 2.1 Utility functions
A function $u(\cdot)\,,\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}$ is a utility function if $u(\mathbf{0})=0$, $u(\cdot)$ is concave, continuous, monotone non-decreasing, and there exists an $\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $u(\mathbf{x})>0$.
We say that a utility function is separable if $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$. We will also distinguish between satiable and non-satiable utilities.
@@
A utility function $u(\cdot)$ is non-satiable if and only if, for any $\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, there exists $\mathbf{x}^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, such that $u(\mathbf{x}^{\prime})>u(\mathbf{x})$. Otherwise, we say the utility function is satiable.
@@
${\mathcal{D}}^{u}(\mathbf{p},b)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\arg\max_{\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\left\{u(\mathbf{x})\,:\,\left\langle\mathbf{p},\mathbf{x}\right\rangle\leq b\right\}\,.$ (1)
Note that if $u$ is a non-satiable utility function, then $\left\langle\mathbf{p},\mathbf{x}\right\rangle=b$ must hold for every $\mathbf{x}\in{\mathcal{D}}^{u}(\mathbf{p},b)$.
${\mathcal{D}}^{u}(\mathbf{p},b)$ is always a non-empty compact set when $\mathbf{p}>0$. However, in cases when some prices may be 0, the supremum of utility over the budget constraint might be $\infty$, resulting in unbounded demand. This may occur even when the supremum is finite, causing technical difficulties. [^5] For example, given $u(x_{1},x_{2})=1-\frac{1}{x_{1}+x_{2}+1}$, $\mathbf{p}=(1,0)$, and $b=1$, $\sup_{\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\left\{u(\mathbf{x})\,:\,\left\langle\mathbf{p},\mathbf{x}\right\rangle\leq b\right\}=1$. This value is approached as $x_{1}=1$ and $x_{2}\rightarrow\infty$, but it is never achieved by any finite bundle $\mathbf{x}$. To circumvent these issues, we adopt the following assumption regarding utility functions $u(\cdot)$:
@@
For any $\mathbf{p}\geq\mathbf{0}$ and $b>0$, if there is a sequence of allocations $\{\mathbf{x}^{(k)}\}_{k=1}^{\infty}$ such that $\left\langle\mathbf{p},\mathbf{x}^{(k)}\right\rangle$ converges to $b$ and $u(\mathbf{x}^{(k)})$ converges to some finite value $M<\infty$, then there exists a finite allocation $\mathbf{x}^{*}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $u(\mathbf{x}^{*})=M$ and $\left\langle\mathbf{p},\mathbf{x}^{*}\right\rangle\leq b$.
This assumption always holds if the set of possible allocations is restricted to a compact domain $Z\subseteq\mathbb{R}_{\geq 0}^{\mathcal{M}}$; note that Nesterov and Shikhman (2018) make such a compactness assumption. For the unrestricted domain $\mathbb{R}_{\geq 0}^{\mathcal{M}}$, Assumption 1 can be shown to hold for arbitrary piecewise linear utility functions, in particular, for generalized network utilities defined below.
@@
$\mathcal{GD}^{u}(\mathbf{p},b)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\arg\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\mathbf{p},\mathbf{y}\right\rangle\,.$ (2)
We define the domain of $\mathcal{GD}^{u}$ with respect to prices as  [^6] An equivalent formulation of the definition is given by: $\operatorname{pdom}(u)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\left\{\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}\penalty 10000\ :\penalty 10000\ \forall b>0,\sup_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}} b\log u(\mathbf{y})-\left\langle\mathbf{p},\mathbf{y}\right\rangle<+\infty\right\}.$ This definition is obtained by substituting the universal quantifier ($\forall$) for the existential quantifier ($\exists$). Lemma A.1 (i) provides further details on this equivalence.
$\operatorname{pdom}(u)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\left\{\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}\penalty 10000\ :\penalty 10000\ \exists b>0,\sup_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\mathbf{p},\mathbf{y}\right\rangle<+\infty\right\}\,.$ (3)
Under Assumption 1, $\mathbf{0}\in\operatorname{pdom}(u)$ if and only if $u(.)$ is satiable. Every strictly positive $\mathbf{p}>\mathbf{0}$ is in $\operatorname{pdom}(u)$, and if $u$ is satiable, then $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$ (see Lemma A.1). In Appendix A, we show the following.
@@
Under Assumption 1, for any $\mathbf{p}\in\operatorname{pdom}(u)$ and for any $b>0$, it holds that $\mathcal{GD}^{u}(\mathbf{p},b)\neq\emptyset$.
Gale demands may behave very differently from the standard demand correspondence. Example E.3 shows the counter-intuitive property increasing the prices $\mathbf{p}$ may actually increase the Gale utility value $\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\mathbf{p},\mathbf{y}\right\rangle$.
Recall the notions of bounded linear, separable piecewise linear (SPLC) and Leontief-free (LF) utilities from the Introduction. We now define the class of generalized network utility functions.
@@
We now formally introduce the generalized flow model. A generalized flow instance is given by a directed graph $G=(V,E)$ with a sink node $t\in V$, a supply set $S\subseteq V\setminus\{t\}$, a capacity vector $\mathbf{h}\in\mathbb{R}_{\geq 0}^{E}$, and positive gain factors $\bm{\gamma}\in\mathbb{R}_{>0}^{E}$. We let $\partial^{\mathrm{in}}(S)$ and $\partial^{\mathrm{out}}(S)$ denote the set of arcs entering and leaving a subset $S\subseteq V$, respectively. For a node $v\in V$, we use $\partial^{\mathrm{in}}(v)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\partial^{\mathrm{in}}(\{v\})$ and $\partial^{\mathrm{out}}(v)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\partial^{\mathrm{out}}(\{v\})$. For a vector $\mathbf{f}\in\mathbb{R}_{\geq 0}^{E}$, for every arc $e=(v,w)$, we interpret $f_{e}$ as the flow entering $e$ at node $v$. The flow gets multiplied by the gain factor $\gamma_{e}$ while traversing the arc, with $\gamma_{e}f_{e}$ amount reaching node $w$. Thus, the net flow at node $v\in V$ is defined as
$\operatorname{net}_{\mathbf{f}}(v)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{e\in\partial^{\mathrm{in}}(v)}\gamma_{e}f_{e}-\sum_{e\in\partial^{\mathrm{out}}(v)}f_{e}\,.$
Given a supply vector $\mathbf{x}\in\mathbb{R}_{\geq 0}^{S}$, we say that $\mathbf{f}\in\mathbb{R}_{\geq 0}^{E}$ is a feasible generalized flow with supply $\mathbf{x}$ and capacity $\mathbf{h}$ if $\mathbf{0}\leq\mathbf{f}\leq\mathbf{h}$, and $\operatorname{net}_{\mathbf{f}}(v)\geq-x_{v}$ for all $v\in S$, and $\operatorname{net}_{\mathbf{f}}(v)\geq 0$ for all $v\in V\setminus(S\cup\{t\})$. The maximum generalized flow value from supply $\mathbf{x}$ is the maximum amount of $\operatorname{net}_{\mathbf{f}}(t)$ among all feasible flows with supply $\mathbf{x}$. It corresponds to the optimum value of the following linear program:
[…]
$\operatorname{net}_{\mathbf{f}}(v)$ $\geq-x_{v}\quad\forall v\in S$
$\operatorname{net}_{\mathbf{f}}(v)$ $\geq 0\quad\forall v\in V\setminus(S\cup\{t\})$
$\mathbf{0}$ $\leq\mathbf{f}\leq\mathbf{h}\,.$
@@
The function $u:\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ is a generalized network utility function if $u(\mathbf{x})$ is the maximum generalized flow value from supply $\mathbf{x}$ in some generalized flow instance where the supply set is $S={\mathcal{M}}$, the maximum flow amount is finite for any supply vector, $u(\mathbf{0})=0$, and there exists $\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ with $u(\mathbf{x})>0$.
The following properties are immediate from the definition and show that these are indeed utility functions in accordance with Definition 2.2.
@@
If $u:\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ is a generalized network utility function, then $u(\mathbf{x})$ is monotone non-decreasing and concave, and satisfies Assumption 1.
@@
An SPLC function can be represented as a generalized network utility function by setting $V={\mathcal{M}}\cup\{t\}$, and for each $j\in{\mathcal{M}}$, adding a capacitated arc $e=(j,t)$ for each linear segment of $v_{j}$, where $\gamma_{e}$ is the marginal utility on this segment and $h_{e}$ is the length of the segment. Similarly, we can represent a Leontief-free utility as a generalized network utility function by setting $V={\mathcal{M}}\cup K\cup\{t\}$; recall $K$ is the set of segments. For each $j\in{\mathcal{M}}$ and $k\in K$, we add an arc $e=(j,k)$ with gain factor $\gamma_{e}=a_{jk}$ and capacity $h_{e}=\infty$. For each segment $k\in K$, we add an arc $e=(k,t)$ with $\gamma_{e}=1$ and $h_{e}=\ell_{k}$.
@@
Assume that agents are competing for access to $m$ machines $\mathcal{M}$} that can produce $\ell$ different products. The machine $j$ can produce $a_{jk}$ units of product $k$ per time unit. The unit revenue of product $k$ is $r_{k}$. There are also limits $t_{jk}$ on the quantity of each product $k$ that the agent can produce on machine $j$, as well a limit $d_{k}$ on the maximum possible amount producible from $k$. Given $x_{j}$ time on machine $j\in{\mathcal{M}}$, the utility of the agent is the maximum possible revenue subject to these constraints. This can be formulated using a two-layer generalized flow network, similar to Leontief-free case described above, but with additional capacity bounds $t_{jk}$. See Figure 2 for an illustration. The model allows for more complex production networks with possibly multiple layers corresponding to raw materials, intermediary and final products.
$x_{1}\bullet$$1\bullet$$x_{2}\bullet$$\vdots$$\vdots$$k\bullet$$\bullet$$x_{i}\bullet$$\vdots$$\vdots$$\ell\bullet$$x_{m}\bullet$$\scriptstyle{(r_{k},d_{k})}$$\scriptstyle{\quad\quad\quad(a_{jk},t_{jk})}\scriptstyle{\vdots}$
Figure 2: Utilities from production.